OpenStudy (aaronandyson):

Please check : A cone of height 8 cm has a curved surface area of 188.4 cm^2.Find its radius. Take pi as 3.14

OpenStudy (aaronandyson):

\[\pi*rl = 188.4\] \[rl = 188.4/\pi = 188.4/3.14 = 60\] Since, \[l = \sqrt{r^2 + h^2}\] \[r\sqrt{r^2 + h^2} = 60\] On squaring both sides \[r^2(r^2 + 64) = 3600\] \[r^4 +64r^2 - 3600 = 0\]

OpenStudy (aaronandyson):

r^2 = 36 cm Therefore r = 6 cm

OpenStudy (baru):

steps are right...